OpenStudy (jjuden):

What is the area of a circle whose radius is 20 meters?

OpenStudy (btaylor):

What is the formula for the area of a circle?

OpenStudy (anonymous):

The area of the circle = Pi x (r^2) Where Pi = 3.14 and r = radius. In this problem, we have Radius(r) = 20 meters => Area of the circle = Pi x (20^2) = 3.14 x 400 = ???

OpenStudy (jjuden):

\[20\pi m ^{2};40\pi m ^{2};200\pi m ^{2};400\pi m ^{2}\] the answers